# LLM Reading Tea Leaves: Automatically Evaluating Topic Models with Large Language Models

### v1

@@ 5.2 Results and Analysis
We assess topic models’ performance based on our evaluation metrics on both 20News and DBpedia. We have the following observations based on our results illustrated in Figure 3: (1) The WALM values of most models on DBpedia show better performance than 20News, which indicates that it is easier for topic models to generate informative topical words for short documents than long documents. (2) The performance indicated by overlap-based metrics (e.g., and ) and embedding-based metrics (e.g., and ) is slightly different. The reason is that embedding-based metrics consider the semantic distance among words, which can be more flexible than the exact match in overlap-based metrics. (3) It can be observed that there is little improvement from recent NTMs over LDA and NVDM in terms of our joint metrics. The potential reason is that most contemporary NTMs primarily focus on enhancing topic coherence while neglecting the generation of documents, thus showing weak performance in generating topical words of documents as indicated by WALM. (4) WALM shows consistent performance ranking across different settings of (e.g., and ) for long documents, which potentially alleviates the effort of extensive model selection in terms of during model comparison.
@@
We illustrate the learning curves of topic models in terms of WALM in Figure 4, which clearly shows how our metrics change along with the learning of a topic model. It demonstrates that our metrics are informative for model selection.
@@
We qualitatively investigate the topical words of documents by topic models at different stages in Table 1, where we randomly sample one document for 20News and DBpedia, respectively. We have the following observations based on our results: (1) The topical words at the beginning phase contain less semantically related words about the documents than those at convergence, which aligns with the learning status (as in Figure 4) indicated by WALM. (2) The topical words of NVDM include more words that reveal the documents’ main messages than LDA, which aligns with the ranking (as in Figure 3) suggested by WALM. (3) The words from LLAMA2 are comparable with those from human annotators in the longer document. For the short document, LLAMA2 focuses on detailed information such as the movie title (e.g., “wrong” and “world”), which human annotators ignore. Further analysis between using LLAMA2 and human judgment to obtain the “true” topical words is described in Sec. 5.4.
@@
We compute Pearson’s correlation coefficients among existing and WALM series metrics, similar to the correlation analysis in previous works such as Doogan and Buntine (2021) and Rahimi et al. (2023). Pearson’s correlation coefficients among the metrics are plotted in a heatmap in Figure 5. Based on the results, we observe that: (1) WALM variants are highly correlated with each other since they originate from the same mechanism. (2) Compared with DC-PPL, which also evaluates the entire model based on documents, WALM show weak correlations, suggesting a new family of evaluation metrics. (3) Compared with other types of evaluations, WALM has moderate correlations with document representation metrics (e.g., KM-Purity, KM-NMI, and ACC), and weak correlations with topic quality metrics (e.g., NPMI and TD). This indicates that our joint evaluation metrics consider both components. Those observations demonstrate the comprehensive nature of WALM, which can be used as a complementary evaluation method to the existing ones.
@@ 5.3 Contextualized Embeddings for WALM
Recall that in Eq. 9 and Eq. 4.4, the cost matrix is constructed using cosine distances between word embeddings. Here, we change our construction of from using static word embeddings from GloVe (Pennington et al., 2014) to the contextualized word embeddings from LLAMA2, considering that the same word may have different semantic meanings in different contexts. We obtain the contextualized embeddings of a word given a document differently in two cases: (1) When the target word appears in the context document, we take the average embeddings of each occurrence as the contextualized embedding. (2) When there is no occurrence of the target word in the given document, we add an auxiliary sentence to the document in the following format:
@@
Since the cost of obtaining contextualized embeddings is high for LLMs, we compute and in a case study, where we test NVDM on 100 documents randomly sampled from the test sets of 20News and DBpedia, respectively. We plot the learning curves on test documents in Figure 6. We observe that using word embeddings or contextualized embeddings in our embedding-based scores exhibits similar trends but with different values on both datasets. Further analysis of both approaches is conducted in Sec. 5.4.
5.4 Comparisons with Human Evaluation
WALM computes the difference between documents’ topical words generated by topic models and an LLM, treating the words from the LLM as the ground truth. Here, we investigate the gap between using LLM and human judgment as the true topical words in WALM. To quantify this gap, we propose the following calculation:
where is the score function we propose in Sec. 4.4 (and the contextualized variants in Sec. 5.3); , , and are the topical words from the topic model, LLM and human annotators, respectively. We empirically observe that the topical words generated by topic models always differ from those identified by humans, thus the denominator in Eq. 12 will not be zero.
Case Study Setup
As human annotation is expensive for large-scale investigation, we randomly sample 200 test documents from 20News and DBpedia as a case study. We ask human annotators to provide keywords that capture the main points of each document. Then, given a trained topic model, we compute the gap between using the words from the LLM and human judgment in our metrics using Eq. 12.
Observations
@@ 6 Conclusion
In this work, we propose WALM (Words Agreement with Language Model), a new evaluation method for topic models, which takes both topic and document representation quality into account jointly. WALM measures the agreement between the topical words generated by topic models and those from the LLM for given documents. We propose different calculations to quantify their agreement, including overlap-based and embedding-based metrics. Through extensive experiments, WALM series metrics show their effectiveness in reflecting topic models’ capability in terms of semantic summary of documents. We show that WALM metrics align with human judgment and can serve as an informative complementary method for topic model evaluation. In the future, we believe that WALM metrics can be further enhanced through improved document summarization by LLMs and word set similarity calculations, and topic models can be further optimized based on WALM series targets.

### v2

@@ 5.2 Results and Analysis
We assess topic models’ performance based on our evaluation metrics on both 20News and DBpedia. We have the following observations based on our results illustrated in Figure 4: (1) The WALM values of most models on DBpedia show better performance than 20News, which indicates that it is easier for topic models to generate informative topical words for short documents than long documents. (2) The performance ranking indicated by overlap-based metrics (e.g., and ) and embedding-based metrics (e.g., and ) is slightly different. The reason is that embedding-based metrics consider the semantic distance among words, which can be more flexible than the exact match in overlap-based metrics. (3) It can be observed that there is little improvement from recent NTMs over LDA and NVDM in terms of our joint metrics. The potential reason is that most contemporary NTMs primarily focus on enhancing topic coherence while neglecting the generation of documents, thus showing weak performance in generating topical words of documents as indicated by WALM. (4) When topic-aware keyword suggestion is applied in WALM (Figure 5), the performance ranking of LDA surpasses that of NVDM as the number of topics increases in the long-document dataset (i.e., 20News). This suggests that LDA benefits more from an increased number of topics when generating topic-aware keywords for documents compared to NVDM.
@@
In Figure 6, we illustrate the learning curves of topic models in terms of WALM, clearly showing how each metric changes throughout the training process. We observe that most topic models improve with training and eventually converge to a stable state. However, NVDM exhibits overfitting in the later stages of training, as indicated by its WALM scores. Additionally, WALM approaches based on keyword suggestions and topic-aware keyword suggestions exhibit slightly different trends in their learning curves. For instance, LDA surpasses NVDM in the later training stages when topic-aware keywords are used. This suggests that NVDM prioritizes document-level generation while LDA shows stronger awareness of collection-level topics.
@@
We qualitatively investigate the topical words of documents by topic models at different stages in Table 1, where we randomly sample one document for 20News and DBpedia, respectively. We have the following observations based on our results: (1) The topical words at the beginning phase contain less semantically related words about the documents than those at convergence, which aligns with the learning status (as in Figure 6) indicated by WALM. (2) The topical words of NVDM include more words that reveal the documents’ main messages than LDA, which aligns with the ranking (as in Figure 5) suggested by WALM. (3) The keywords generated by the LLM are similar to those provided by human annotators for the example documents. (4) By using topic-aware keywords suggestion in WALM, the LLM tends to provide keywords that convey the high-level concepts of the topics. For instance, “troubleshooting” is identified for the first example document, and “entertainment” for the second, which offers higher-level information from topics besides individual document.
@@
We compute Pearson’s correlation coefficients among existing and WALM series metrics, similar to the correlation analysis in previous works such as Doogan and Buntine (2021) and Rahimi et al. (2024). Pearson’s correlation coefficients among the metrics are plotted in a heatmap in Figure 7. Based on the results, we observe that: (1) WALM variants are highly correlated with each other since they originate from the same mechanism. (2) Compared with perplexity, which also evaluates the entire model based on documents, WALM show weak correlations, suggesting a new family of evaluation metrics. (3) Compared with other types of evaluations, WALM has moderate correlations with document representation metrics (e.g., KM-Purity, KM-NMI, and ACC), and weak correlations with topic quality metrics (e.g., NPMI and TD). This indicates that our joint evaluation metrics take both components into account without relying solely on either one. These observations suggest that WALM can serve as a complementary evaluation method to existing approaches.
@@ 5.3 Contextualized Embeddings for WALM
Recall that in Eq. 9 and Eq. 4.4, the cost matrix is constructed using cosine distances between word embeddings. Here, we change our construction of from using static word embeddings from GloVe (Pennington et al., 2014) to the contextualized word embeddings from the LLM, considering that the same word may have different semantic meanings in different contexts. We obtain the contextualized embeddings of a word given a document differently in two cases: (1) When the target word appears in the context document, we take the average embeddings of each occurrence as the contextualized embedding. (2) When there is no occurrence of the target word in the given document, we add an auxiliary sentence to the document in the following format:
@@
Since the cost of obtaining contextualized embeddings is high for LLMs, we compute and in a case study, where we test NVDM on 100 documents randomly sampled from the test sets of 20News and DBpedia, respectively. We plot the learning curves on test documents in Figure 8. We observe that using word embeddings or contextualized embeddings in our embedding-based scores exhibits similar trends but with different values on both datasets.
5.4 Sensitivity Study
Here, we examine two factors that can influence the WALM scores: the number of keywords generated by the LLM and the choice of the LLM. To investigate the effect of the number of keywords, we vary the number from 3 to 10 and plot the performance ranking of topic models in Figure 9 (top row). We observe that, although the values of WALM metrics can vary with different numbers of keywords, the overall performance ranking of the topic models remains largely unaffected by these changes, especially for the overlap-based metrics. To investigate the effect of LLMs, we use different latest LLMs for keyword generation apart from LLAMA3-8B-Instruct, including Mistral-7B-Instruct-v0.3 (Jiang et al., 2023), Phi-3-Mini-128K-Instruct (Abdin et al., 2024) and Yi-1.5-9B-Chat (Young et al., 2024). From the results illustrated in Figure 9 (bottom row), we observe that overlap-based metrics show minimal variation with different choices of LLMs, and the performance ranking of the topic models is unaffected in most cases. These observations suggest that the overlap-based metrics are less sensitive to the number of words and the choice of LLMs.
5.5 Comparisons with Human Annotation
Evaluation Gap with Human Annotation
WALM computes the difference between documents’ topical words generated by topic models and an LLM, treating the words from the LLM as the ground truth. Here, we investigate the gap between using LLM and human judgment as the true topical words in WALM. To quantify this gap, we use the following calculation:
where is the WLAM scores we propose in Sec. 4.4 (and the contextualized variants in Sec. 5.3). Intuitively, the gap function measures the difference between using ground truth (e.g., keywords) from the LLM and those from human annotators. We empirically observe that topical words from topic models consistently differ from those identified by humans, so the denominator in Eq. 12 will not be zero.
As human annotation is expensive for large-scale investigation, we randomly sample 200 test documents from 20News and DBpedia as a case study. We engaged three English speakers as annotators, trained with a few examples, to provide keywords that capture the main points of each document. Then, given a trained topic model, we compute the gap between using the words from the LLM and human in our metrics using Eq. 12.
@@ 5.5 Comparisons with Human Annotation
5-word suggestion | 0.55 | 0.50 | 0.57 | 0.63 |
10-word suggestion | 0.52 | 0.58 | 0.56 | 0.68 |
Correlation with Human Annotation
We use an existing annotated dataset, 500N-KPCrowd (Marujo et al., 2012) for the keyphrase extraction task (Hasan and Ng, 2014), where each test document is paired with labeled keywords. We run LDA on the training documents and infer the topical words for the test documents, then compute the Pearson’s correlation coefficient between the WALM scores using the LLM-generated keywords and the test labels as the ground truth. The results are illustrated in Table. 2. We observe that (1) using keywords from the LLM in WALM scores correlates with using the labeled keyphrases, and (2) the correlation can potentially improve when more keywords are included in the LLM’s suggestions.
@@ 6 Conclusion
In this work, we propose WALM for topic model evaluation, which takes both topic and document representation quality into account jointly. WALM measures the agreement between the topical words generated by topic models and those from the LLM for given documents. The topical words from the LLM are obtained through keyword prompting or topic-aware keyword prompting, with the latter tending to capture higher-level information. To quantify the agreement between word sets, we propose different calculations, including overlap-based and embedding-based metrics. Our experiments demonstrate that the WALM series effectively reflect the capability of topic models to provide semantic summaries of documents. We show that WALM metrics align with human judgment and can serve as an informative complementary method for topic model evaluation. We suggest that overlap-based metrics demonstrate better sensitivity handling, while embedding-based metrics show a smaller evaluation gap. A potential risk of using WALM is that models chasing this metric only may be affected by the bias of LLMs. To mitigate the risk, we suggest using WALM with other metrics together.
Acknowledgments
We thank the anonymous reviewers and the action editor, Michael Elhadad, for their valuable feedback, which has significantly strengthened this work.